OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (amistre64):

in other words, give them an answer which constitutes what you have gained thru the exercies

OpenStudy (amistre64):

how would you go about solving these? other than relying upon other people to do it for you is the question they are asking :)

OpenStudy (anonymous):

i have to add that this "complete sentence" motif is getting silly. explain in complete sentences how you walk down a flight of stairs...

OpenStudy (amistre64):

first i say a prayer, then I jump :)

OpenStudy (anonymous):

use matrix,.

OpenStudy (amistre64):

Without knowing which methods you have and have not been taught, our answers will not be astute

OpenStudy (amistre64):

matrix is only really viable by machine; otherwise it just takes too long in my opinion

OpenStudy (amistre64):

id use descartes method and graph circles and lines intersecting at odd angles to imagined dots and add up the circles

OpenStudy (anonymous):

i do not rely on anyone to solve this for me i just do not know where to start from

OpenStudy (anonymous):

although it's an exhaustful process, but there is an assurance at the end of your computation, your answer is reliable. .

OpenStudy (amistre64):

you should start by choosing a method :)

OpenStudy (amistre64):

substitution is the most reliable method to me as long as your mathing is up to par

OpenStudy (amistre64):

elimination is quicker but prone to violent fits of fury and rage

OpenStudy (amistre64):

matrix is a combonation of the 2

OpenStudy (amistre64):

graphing is also a method but is the least reliable

OpenStudy (anonymous):

thank you so much i have a better understanding of it now thanks for all your help

OpenStudy (amistre64):

yw, and good luck with it ;)

OpenStudy (amistre64):

just curious about something .... 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11 2x + y + z +7=0 x – 3y + 4z +14=0 x – 2y – 3z +11=0 2x + y + z +7 (=) x – 3y + 4z +14 (=) x – 2y – 3z +11 -x + 2y + 3z -11 -x+ 2y + 3z -11 -x+ 2y + 3z -11 ------------------------------------------------ x + 3y +4z -4 (=) -y +7z +3 x + 3y +4z -4 (=) -y+7z+3 +y -7z -3 +y -7z -3 -------------------------- x +4y -3z -7 = 0 hmmm, i got no idea what that could tell us about a solution tho

OpenStudy (amistre64):

2 planes intersect in a line; <2,1,1> <1,-3,4> x 2 1 x =4--3 = 7 y 1 -3 -y =(8-1) = -7 ; <1,-1,-1> is our direction vector then z 1 4 z =-6-1 = -7 and a point that matches on both planes is: 3y +3z = –21 .... y-5=-7; y=-2 -3y + 4z = –14 --------------- 7z = -35 ; z=-5 (0,-2,-5) our line is then defined as: x= 0+t y=-2-t z=-5-t insert these xyz into the third equation and solve for t x – 2y – 3z = –11 t +4+2t +15 +3t = -11 6t = -30 ... t=-5 ................................ plug this into our line to see what point we get :) x= 0+-5 = -5 y=-2--5 = 7 z=-5--5 = 10 (-5,7,10) might be it if i played me cards roght :)

OpenStudy (amistre64):

z = 0 ... y = 3 i forgot how to add

OpenStudy (amistre64):

2x + y + z = –7 -5 3 0 ------------ -10+3 good x – 3y + 4z = –14 -5 3 0 ----------- -5-9 good x – 2y – 3z = –11 -5 3 0 ---------- -5-6 good

OpenStudy (amistre64):

anyways :) as you can see, there are quite a few ways to solve this; some more complicated than others ;)